OpenStudy (anonymous):

if x^2+xy=10 then whenx=2, dy/dx=?

OpenStudy (anonymous):

Use implicit differentiation. In my notation , y' is dy/dx x^2 + xy = 10 2x + y + xy' = 0 What is y when x is 2? Go back to the original and plug in 2 for x. You get 2^2 + 2y = 10, and you get y = 3. So x =2, y=3: plug that into 2x + y + xy' = 0...solve for y' and that's your answer.